OpenStudy (anonymous):

Convert the following fraction to a decimal. -5/8 A.-0.658 B.0.65 C.-0.625 D.0.855

OpenStudy (alexandervonhumboldt2):

what do you think?

OpenStudy (anonymous):

idk I learned how to convert fractions into decimals but the next day forgot

OpenStudy (alexandervonhumboldt2):

ok so 1000/8=125. so now you have to multiply the -5/8 * 125/125=(-5*125)/(8*125)=???

OpenStudy (alexandervonhumboldt2):

no

OpenStudy (anonymous):

Still need help?